Calculate the kinetic energy of a neutron whose momentum is 200 MeV/c


1
Expert's answer
2021-04-12T06:58:03-0400

Answer

The kinetic energy is given by

"K=\frac{P^2}{2m_n}\\=\frac{(200MeV/c)^2}{2\times1.67\times10^{-27}}\\=\frac{(40000\times1.6\times10^{-13})}{2\times1.67\times10^{-27}}"

"=1.92\times10^{18}J"